What is the principal:
Rate at 12%
Time is 2 months
Simple interest is $960

Answer by Theo(13342)

your percent rate is assumed to be your annual percent rate.
the monthly rate percent is 12% / 12 = 1 percent per month.
the monthly rate is the percent divided by 100 = 1/100 = .01

your time is 2 months.

your simple interest is $960.

formula for simple interest is i = p * r * n

i is the interest
p is the principal
r is the interest rate per time period
n is the number of time periods.

in your problem, you have:

i = 960
p = what you want to find.
r = .01 per month.
n = 2 months

your formula becomes:

960 = p * .01 * 2

solve for p to get p = 960 / (.01 * 2) = 48,000

your principal is 48,000
your interest rate is .01 per month.
your number of months is 2
your interest is 48,000 * .01 * 2 = 960